OpenStudy (anonymous):

A survey of 1,000 men and women asked, "Do you earn over $50,000 per year?" The table below shows the responses for males and females.

OpenStudy (anonymous):

Male Female Total Income over $50,000 485 385 870 Income below $50,000 65 65 130 Total 550 450 1,000

OpenStudy (anonymous):

Based on these data, are "being male" and "earning over $50,000" independent events? Yes, P(being male | the person earns over $50,000) ≠ P(being male) Yes, P(being male | the person earns over $50,000) = P(being male) No, P(being male | the person earns over $50,000) ≠ P(being male) No, P(being male | the person earns over $50,000) = P(being male)

OpenStudy (anonymous):

Independent events: P(A and B) = P(A)P(B) P(male, >50k) = 485/1000 = .485 P(male) = 550/1000 = .550 P(>50k) = 870/1000 = .870 P(male)P(>50k) = .4785 not 0.485 thus not independent The options seem to use another criterion, however.

OpenStudy (anonymous):

Yea I kind of do not really get what the question is asking... I am super confused.

OpenStudy (anonymous):

I guess if independent then P(being male and >50k/yr) = P(being male) because making >5ok would not make you more or less likely to be a male.

OpenStudy (anonymous):

Yes, I was going to ask you if you thought that it would be D... I will just go with that. It makes the most sense I think.

OpenStudy (anonymous):

Thank you for your help (:

OpenStudy (anonymous):

We were not correct... I think it was actually C lol